How many grams would 132 moles of C3H5O3 weigh?
1
Expert's answer
2018-05-11T04:25:00-0400
n(C3H5O3) = 132 moles

M(C3H5O3) = 12*3 + 1*5 + 16*3 = 89 g/mol

m(C3H5O3) = n*M = 132mol*89g/mol = 11748 g